# Gravity from entropy

### v4

@@ Abstract
Gravity is derived from an entropic action coupling matter fields with geometry. The fundamental idea is to relate the metric of Lorentzian spacetime to a density matrix and to describe the matter fields topologically, according to a Dirac-Kähler description, as the direct sum of a zero-form, a one-form and a two-form. While the geometry of spacetime is defined by its metric, the matter fields can be used to define an alternative metric, the metric induced by the matter fields, which geometrically describes the interplay between spacetime and matter. The proposed entropic action is the quantum relative entropy between the metric of spacetime and the metric induced by the matter fields. The modified Einstein equations obtained from this action reduce to the Einstein equations with zero cosmological constant in the regime of low coupling. By introducing the G-field, which acts as a set of Lagrangian multipliers, the proposed entropic action reduces to a dressed Einstein-Hilbert action with an emergent small and positive cosmological constant only dependent on the G-field. The obtained equations of modified gravity remains second order in the metric and in the G-field. A canonical quantization of this field theory could bring new insights into quantum gravity while further research might clarify the role that the G-field could have for dark matter.
@@ I Introduction
The relation between general relativity, statistical mechanics and information theory is a central research topic in theoretical physics. The interest in the subject has its roots in the discovery that black holes have an entropy [1, 2] and emit Hawking radiation [3]. Recently, important results have been obtained relating information theory, entanglement entropy [4, 5, 6] and gravity [7, 8, 9, 10, 11, 12, 13] involving the holographic principle [14, 15, 16], the entanglement properties of quantum field theory and the theory of von Neumann algebras [17, 18, 19].
@@
In this work, a continuum modified theory of gravity based on a statistical mechanics action is considered. This theory treats the metric at each point of spacetime as a density matrix, this central idea being inspired by the use of von Neumann algebras in explaining entanglement in field theories [17, 18]. While the geometry of spacetime is defined through its associated metric, the interplay between matter fields and geometry is captured by the metric induced by the matter fields which describes how the matter fields effectively curves spacetime. Embracing a statistical mechanics approach to gravity, this work proposes an action based on the quantum entropy between the metric and the metric induced by the matter fields. Thus the proposed action is meant to fully describe how matter curves geometry and how geometry affects the matter fields. A crucial aspect of the presented theory is the adoption of a topological (Dirac-Kähler like [34, 35]) description of bosonic matter fields. Note that the extension of Dirac-Kähler and staggered fermions formalism to bosonic particles is gaining increasing interest in lattice gauge theory [36, 37] and in network theory as well [38]. These bosonic matter fields are described as the direct sum of a 00-form, a $1$-form, and a $2$-form defined on the Riemannian manifold describing spacetime. Moreover the metric induced by the topological matter fields is expressed in terms of the Hodge-Dirac operator [39]. From this statistical mechanics approach to gravity we derive the modified Einstein equations by introducing an auxiliary field associated to gravity which we call G-field. The introduction of this new field is justified as it acts as a set of Lagragian multipliers enforcing linear constraints on metric induced by the matter fields. In this way the G-field extends the popular use of Lagrangian multipliers in $f(R)$ theories [23, 21]. Given the particular entropic structure of the action, the modified Einstein equations take a very simplified expression. The gravitational part of the action takes the form of a dressed Einstein-Hilbert action in which we observe an emergent positive cosmological constant that depends exclusively on the G-field.
@@
To keep the discussion concise the focus is here exclusively on scalar (bosonic) matter fields, and their topological generalizations while in Ref. [38] the theory covers also fermionic matter fields and gauge fields. Further extensions of the proposed local framework to Dirac and gauge fields [38, 42, 43] in the continuum or in the discrete setting are left for future investigations.
@@ II.1 Eigenvalues and logarithm of rank 2-tensors
$[\ln(\hat{\bf G}^{-1}]^{\mu\nu}=-[V^{(\lambda)}]^{\mu}[V^{(\lambda)}]^{\nu}\ln(\lambda).$ (5)
@@
In this first section we are interested exclusively on rank-two tensors that are metrics between $1$-forms. In the subsequent paragraphs and in Appendix C we will extended the notion of eigenvalues also to metric matrices between two-forms represented by rank-4 tensors. Such an approach will be shown in Appendix C to be general and applicable to metric tensors between two $n$-forms with $n$ of any order.
@@ II.2 Warm-up scenario: Entropic action
Before we develop our theory, let us consider an instructive warm-up scenario that will help justify the theory that we will present in the following. Having defined the logarithm of positively defined rank-2 tensors, we are now in the position to define their Lorentz invariant quantum entropy $H$ for metric associated to $1$-forms. This is inspired by the expression of the Von Neumann entropy albeit we do not require the tensor to have trace one at every point in spacetime. We consider positive definite, invertible metric for $1$-forms, represented by the covariant tensor $\hat{\bf G}$ of rank two. We define the entropy $H$ of $\hat{\bf G}$ as
@@
Note that the problem of defining eigenvalues and entropy of tensors is a topic of intense research and similar definitions are been provided for instance in the theory of elasticity [44] and in applied tensor analysis [45] (although not defined in a Lorentz invariant way). Recalling that the metric $g$ for $1$-forms has all the eigenvalues equal to one, it follows that the entropy of of the metric $g$ is null, i.e.
@@
where $|-g|$ indicates the absolute value of the determinant of $g$.
A series of observation are here in place to justify our interpretation of the metrics matrices as density matrices. These considerations are important for developments that might lead to a full quantum theory upon canonical quantization. As remarked previously we treat the metric matrices as density matrices with some important differences, including the fact that we require that they are invertible and we do not require that they have unitary trace. The requirement of treating exclusively invertible metrics $g$ and ${\bf G}$ is dictated by our desire to define the entropy in a Lorentz invariant way. In fact here we desire to treat the metric ${\bf G}$ on the same footing of its inverse ${\bf G}^{-1}$. It might be argued that also our relaxation of the requirement to have density matrices of trace one, has similar roots. In fact if we require ${\bf G}$ to have unitary trace its inverse in general will not have a unitary trace.
The relaxation of the condition to have unitary trace density matrices is related to the theory of quantum entanglement and the use of von Neumann algebras [17, 18]. Physically, we assume that the tensor $\hat{\bf G}$ that is fully characterizing the geometry of the entire manifold, is defined at each point of $p$ of the manifold $\mathcal{K}$. It might be tempting to describe this geometry as an uncountable tensor product of the metrics defined at each point $p$. Thus, unless we integrate over a suitable measure, the trace associated to the entire metric is not defined. However, at each point $p$ the metric $\hat{\bf G}$ is associated to a finite trace, and to a finite entropy $H$ which depends on $p$. This interpretation of the metric matrices $\tilde{\bf G}$ is in line with the physical understanding of von Neumann algebras [18, 17], although the establishment of a mathematical rigorous relation of the present approach with von Neumann algebras is beyond the scope of this work and will be the subject of subsequent publications.
The interpretation of the metric matrix $g$ as a density operator, also requires addressing which is the underlying quantum statistical ensemble in this physical interpretation. Although it is not our goal here to develop a full quantum theory, an indication of the underlying quantum interpretation might be provided by the introduction of quantum forms or q-forms (see Appendix A). These q-forms might represent the quantum state of geometry and correspond to a geometric version of q-bits. Thus the metric $\hat{\bf G}$ can be interpreted as a “renormalizable density matrix” acting on the Hilbert space of q-forms.
@@ II.3 Warm-up scenario: scalar matter fields
We now apply this warm-up scenario in presence of scalar matter fields. We consider the complex valued scalar matter field $\phi({\bf x})\in\mathbb{C}$ with ${\bf x}\in\mathcal{M}$ and we consider the $d$-dimensional manifold immersed in $\mathcal{K}\otimes\mathbb{C}$ defined by the points $({\bf x},\phi({\bf x}))$. The metric ${\bf G}$ induced on $\mathcal{M}$ by this matter fields is given by
@@
In the limit $\alpha w\to 0$ we have $h(w)\to\alpha$ and the equation for the scalar field reduces to the simple massless Klein-Gordon equation. As we have already discussed previously, one might wonder whether our choice to take directly the metric matrix and interpret it as a “renormalizable density matrix” is the most proper one, and whether it would be better to instead renormalize directly $g$ and ${\bf G}$ with their trace. As an additional possible answer to this important question, in Appendix B we investigate what would be the corresponding equation of motion in this warm-up scenario. The result of this simple exercise points out that if we renormalize $g$ and ${\bf G}$ with their trace at each point $p$ of the manifold, we will not obtain the Klein-Gordon equation in the linear approximation, in part justifying our approach inspired from von Neumann algebras.
By putting equal to zero the variation of the action $\delta\mathcal{S}=0$ with respect to the metric $g$, we get
@@ III.1 Topological matter fields and their associated metrics
The considered metric matrix $\tilde{g}$ is defined as the direct sum of the metric matrix for 00-forms (the identity), the metric matrix $g$ among $1$-forms introduced previously, and the metric $g_{(2)}$ for $2$-forms given by
$g_{(2)}=g_{\mu\nu}g_{\rho\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma}),$ (26)
@@
The scalar product among topological fields is defined as
@@
where $\omega^{\mu}\omega_{\mu}=\omega_{\nu}g^{\nu\mu}\omega_{\mu}$, while the outer product is given by
@@
In our interpretation of the metric as a quantum density matrix, this would correspond to density matrix corresponding to a pure state $D\ket{\Phi}$. However, if the induced metric $\tilde{\bf G}$ is interpreted as a density matrix, it is natural to add further terms in the metric induced by the topological matter field $\tilde{\bf G}$. These terms will depend on the topological field and the geometry of the space which will allow us to describe mixed states. First, we introduce in $\tilde{\bf G}$ a term proportional to |$\ket{\Phi}\bra{\Phi}$ |. Specifically we introduce a term $(m^{2}+\xi R)\ket{\Phi}\bra{\Phi}$ | where $R$ is the Ricci scalar and $\xi$ might include the case of conformal coupling $\xi=1/6$. This term can be also interpreted as a projector. Secondly, we introduce a term depending explicitly on the curvature of the manifold. Since $\tilde{\bf G}$ involves the metric for 00-forms, $1$-forms and $2$-forms on equal footing, it is natural to consider a further term involving $\tilde{\bm{\mathcal{R}}}$ given by the direct sum of the Ricci scalar $R$, the Ricci tensor of elements $R_{\mu\nu}$ and the Riemann tensor of elements $R_{\mu\nu\rho\sigma}$, i.e.
$\tilde{\bm{\mathcal{R}}}$ $=$ $R\oplus(R_{\mu\nu}dx^{\mu}\otimes dx^{\nu}\Big{)}$ (36)
@@
Possibly this approach could be extended to include higher-forms. However, for simplicity, we consider here only topological matter fields formed by the direct sum between a 0-form, a 1-form and a 2-form as this the minimal choice that will allow us to include in the action the Ricci scalar, the Ricci and the Riemann tensor explicitly.
@@ III.2 Entropic topological and geometrical action
We propose a statistical mechanics action formulated in terms of the quantum relative entropy between the metric $\tilde{g}$ and the metric $\tilde{\bf G}$ induced by the topological matter fields. In order to define our action we need to extend the notion of eigenvalues to metric tensors between $2$-forms. This will allow to define the entropy and the quantum relative entropy in this novel framework. For a detailed discussion of this mathematical background see Appendix C that constitutes also the reference for our notation. First of all we observe that the entropy associated to the metric $\tilde{g}$ remains zero. Indeed we define $\tilde{H}$ as
@@
$=$ $1\ln 1+\mbox{Tr}g\ln g^{-1}=\mbox{Tr}g_{(2)}\ln g_{(2)}^{-1}=0,$
where we have used Eq.(9) and the identity derived in the Appendix C (in Eq.(119)),
@@
Since the metric matrices $g_{(n)}$ have all their eigenvalues equal to one, using the notation developed in Appendix C we can express this Lagrangian as well as
@@
The resulting statistical mechanics action $\mathcal{S}$ associated to the Lagrangian $\mathcal{L}$ is given by
@@
$\mathcal{L}$ $=$ $3{\beta}R-\alpha\bra{\Phi}{D}^{2}\ket{\Phi}$ (46)
@@
$\bra{\Phi}{D}^{2}\ket{\Phi}$ $=$ $|\nabla\phi|^{2}+|\nabla^{\mu}\omega_{\mu}|^{2}+\Big{|}\nabla^{\sigma}\bar{\omega}_{\sigma\mu}\Big{|}^{2}$ (47)

### v5

@@ Abstract
Gravity is derived from an entropic action coupling matter fields with geometry. The fundamental idea is to relate the metric of Lorentzian spacetime to a quantum operator, playing the role of an renormalizable effective density matrix and to describe the matter fields topologically, according to a Dirac-Kähler formalism, as the direct sum of a zero-form, a one-form and a two-form. While the geometry of spacetime is defined by its metric, the matter fields can be used to define an alternative metric, the metric induced by the matter fields, which geometrically describes the interplay between spacetime and matter. The proposed entropic action is the quantum relative entropy between the metric of spacetime and the metric induced by the matter fields. The modified Einstein equations obtained from this action reduce to the Einstein equations with zero cosmological constant in the regime of low coupling. By introducing the G-field, which acts as a set of Lagrangian multipliers, the proposed entropic action reduces to a dressed Einstein-Hilbert action with an emergent small and positive cosmological constant only dependent on the G-field. The obtained equations of modified gravity remain second order in the metric and in the G-field. A canonical quantization of this field theory could bring new insights into quantum gravity while further research might clarify the role that the G-field could have for dark matter.
@@ I Introduction
The relation between general relativity, statistical mechanics and information theory is a central research topic in theoretical physics. The interest in the subject has its roots in the discovery that black holes have an entropy [1, 2] and emit Hawking radiation [3]. Recently, important results have been obtained relating information theory, entanglement entropy [4, 5, 6] and gravity [7, 8, 9, 10, 11, 12, 13] involving the holographic principle [14, 15, 16], the entanglement properties of quantum field theory and the theory of von Neumann algebras [17, 18].
@@
In this work, a continuum modified theory of gravity based on a statistical mechanics action is considered. This theory treats the metric at each point of spacetime as a renomalizable density matrix, or more precisely a local quantum operator. This central idea is relating geometry with the mathematical foundation of quantum field theory [33] and is inspired by the used of von Neumann algebras in explaining entanglement in field theories [17, 18] and quantum gravity  [34, 35, 36]. While the geometry of spacetime is defined through its associated metric, the interplay between matter fields and geometry is captured by the metric induced by the matter fields which describes how the matter fields effectively curves spacetime. Embracing a statistical mechanics approach to gravity, this work interprets these metric tensors as quantum operators and postulates an action for gravity given by the quantum relative entropy between the metric of the manifold and the metric induced by the matter fields. From the mathematical point of view, the quantum relative entropy proposed in this work is strictly related to the Araki quantum relative entropy for von Neumann algebras [37, 38, 17, 39]. From the physics point of view, the proposed action fully describes how matter curves geometry and how geometry affects the matter fields.
A crucial aspect of the proposed theory is the adoption of a topological (Dirac-Kähler like [40, 41]) description of bosonic matter fields. Note that the extension of Dirac-Kähler and staggered fermions formalism to bosonic particles is gaining increasing interest in lattice gauge theory [42, 43] and in network theory as well [44]. These bosonic matter fields are described as the direct sum of a 00-form, a $1$-form, and a $2$-form defined on the Riemannian manifold describing spacetime. Moreover the metric induced by the topological matter fields is expressed in terms of the Hodge-Dirac operator [45]. From this statistical mechanics approach to gravity we derive the modified Einstein equations by introducing an auxiliary field associated to gravity which we call G-field. The introduction of this new field is justified as it acts as a set of Lagragian multipliers enforcing linear constraints on metric induced by the matter fields. In this way the G-field extends the popular use of the Legendre transformation in $f(R)$ theories [22, 20]. Given the particular entropic structure of the action, the modified Einstein equations take a very simplified expression. The gravitational part of the action takes the form of a dressed Einstein-Hilbert action in which we observe an emergent positive cosmological constant that depends exclusively on the G-field.
@@
To keep the discussion concise the focus is here mostly on scalar (bosonic) matter fields, and their topological generalizations with a brief mention in Appendix A to the natural extension of this framework to Abelian gauge fields, while in Ref. [44] the theory covers also fermionic matter fields. Further extensions of the proposed local framework to Dirac and non-Abelian gauge fields [44, 48, 49] in the continuum or in the discrete setting are left for future investigations.
@@ II.1 Eigenvalues and logarithm of rank 2-tensors
$[\ln(\hat{\bf G}^{-1})]^{\mu\nu}=-[V^{(\lambda)}]^{\mu}[V^{(\lambda)}]^{\nu}\ln(\lambda).$ (5)
@@
In this first section we are interested exclusively on rank-two tensors that are metrics between vectors (and 1-forms). In the subsequent paragraphs and in Appendix B we will extended the notion of eigenvalues also to metric matrices between bivectors (and 2-forms) represented by rank-4 tensors. Such an approach will be shown in Appendix B to be general and applicable to metric tensors between two $n$-vectors (an $n$-forms) with $n$ of any order.
Covariant Metrics Elements Definition
$g$ $g_{\mu\nu}$ default covariant rank-2 metric tensor between vectors associated to the manifold $\mathcal{K}$
${\bf G}$ $G_{\mu\nu}$ covariant rank-2 metric between vectors induced by the matter fields
$\hat{\bf G}$ $\hat{G}_{\mu\nu}$ general covariant rank-2 metric tensor including both $g$ and ${\bf G}$
Table 1: Notation used to indicate the different metrics in the warm-up-scenario
@@ II.2 Warm-up scenario: Entropic action
Before we develop our theory, let us consider an instructive warm-up scenario that will help justify the theory that we will present in the following. Having defined the logarithm of positively defined rank-2 tensors, we are now in the position to define their Lorentz invariant quantum entropy $H$ for metric associated to $1$-forms. This is inspired by the expression of the Von Neumann entropy albeit we do not require the tensor to have trace one at every point in spacetime. Thus strictly speaking we interpret the generic metric tensor as a quantum operator [33], which has the physical interpretation of a renormalizable effective density matrix [18]. We consider positive definite, invertible metric for $1$-forms, represented by the covariant tensor $\hat{\bf G}$ of rank two. We define the entropy $H$ of $\hat{\bf G}$ as
@@
Note that the problem of defining eigenvalues and entropy of tensors is a topic of intense research and similar definitions are been provided for instance in the theory of elasticity [50] and in applied tensor analysis [51] (although not defined in a Lorentz invariant way). In the present warm-up scenario we will define the entropy and the quantum relation entropy of different metrics taking the form of rank-2 tensors. As a reference for the notation used, the reader can refer to Table 1. Recalling that the metric $g$ has all the eigenvalues equal to one, it follows that the entropy of of the metric $g$ is null, i.e.
@@
where $|-g|$ indicates the absolute value of the determinant of $g$ and $\ell_{P}$ indicates the Planck length.
We note that the quantum relative entropy [37, 38, 17] is a central quantity in quantum information [39], in the theory of local quantum operators [33] and the mathematical foundations of quantum gravity  [34, 35, 36].
Although we are not aware of previous interpretations of the metrics as quantum operators, it is well known that the quantum relative entropy can also be defined among quantum operators that generalize density matrices. In particular the Araki quantum relative entropy [37, 38]. is defined between quantum operators, that like our metric matrices admit a finite trace at each point of the manifold but this trace might be not unitary and be a function of the considered point of the manifold.
While we leave the discussion of the relation of our entropy action to the Araki quantum relative entropy to Appendix C, a series of physical observations are here in place to justify our interpretation of the metrics matrices as quantum operators. As remarked previously we treat the metric tensors as quantum operators or effective density matrices. The main differences between the metric tensor and the density operators include the fact that we require that the metric matrices are invertible and we do not require that they have unitary trace. The requirement of treating exclusively invertible metrics $g$ and ${\bf G}$ is dictated by our desire to define the entropy in a Lorentz invariant way. In fact here we desire to treat the metrics $g$ and ${\bf G}$ on the same footing of their inverse $g^{-1}$ and ${\bf G}^{-1}$. It might be argued that also our relaxation of the requirement to have density matrices of trace one, has similar roots. In fact if we require $g$ and ${\bf G}$ to have unitary trace at each point of the manifold their inverse in general will not have a unitary trace.
@@ II.3 Warm-up scenario: scalar matter fields
We now apply this warm-up scenario in presence of scalar matter fields. We consider the complex valued scalar matter field $\phi({\bf x})\in\mathbb{C}$ with ${\bf x}\in{\mathcal{K}}$ and we consider the $d$-dimensional manifold immersed in $\mathcal{K}\otimes\mathbb{C}$ defined by the points $({\bf x},\phi({\bf x}))$. The metric ${\bf G}$ induced on $\mathcal{K}$ by this matter fields is given by
@@
In the limit $\alpha w\to 0$ we have $h(w)\to\alpha$ and the equation for the scalar field reduces to the simple massless Klein-Gordon equation. By putting equal to zero the variation of the action $\delta\mathcal{S}=0$ with respect to the metric $g$, we get
@@ III The entropic theory of matter fields coupled to geometry
Covariant Topological Metrics Interpretation
$\tilde{g}=1\oplus g_{\mu\nu}dx^{\nu}\otimes dx^{\nu}\oplus{[g_{(2)}]}_{\mu\nu\rho\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma})$ default covariant topological metric between
the topological fields associated to the manifold $\mathcal{K}$
$\tilde{\bf G}=G_{(0)}+[G_{(1)}]_{\mu\nu}dx^{\mu}\otimes dx^{\nu}\oplus[G_{[2]}]_{\mu\nu\rho\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma})$ covariant topological metric induced by the matter fields
Table 2: Covariant topological metrics used in the general scenario.
Covariant metric tensors Elements Intepretation
$1$ $1$ default metric tensor between scalars
$g$ $g_{\mu\nu}$ default metric tensor between vectors
$g_{(2)}$ ${[g_{(2)}]}_{\mu\nu\rho\sigma}$ default metric tensor between bivectors
${G}_{(0)}$ ${G}_{(0)}$ metric tensor between scalars induced by the matter fields
${\bf G}_{(1)}$ ${[{\bf G}_{(1)}]}_{\mu\nu}$ metric tensor between vectors induced by the matter fields
${\bf G}_{(2)}$ ${[{\bf G}_{(2)}]}_{\mu\nu\rho\sigma}$ metric tensor between bivectors induced by the matter fields
Table 3: Covariant metric tensors between $n$-vectors that are included in the covariant topological metrics.
@@ III.1 Topological matter fields and their associated metrics
The considered covariant metric $\tilde{g}$ is defined as the direct sum of the metric among scalars (the identity), the metric $g$ among vectors introduced previously, and the metric $g_{(2)}$ among bivectors given by
$g_{(2)}=g_{\mu\rho}g_{\nu\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma}),$ (26)
@@
The local scalar product among topological fields is defined as
@@
where $\omega^{\mu}\omega_{\mu}=\omega_{\nu}g^{\nu\mu}\omega_{\mu}$, while the local outer product is given by
@@
Thus in this general scenario we consider the two covariant topological metrics $\tilde{g}$ and $\tilde{\bf G}$ each given by the direct sum of metrics between scalars, vectors and bivectors. As a reference on our notation for this general scenario we refer the reader to the Table 2 and the Table 3.
In our interpretation of the metric as a quantum density matrix, this would correspond to density matrix corresponding to a pure state $D\ket{\Phi}$. However, if the induced metric $\tilde{\bf G}$ is interpreted as a density matrix, it is natural to add further terms in the metric induced by the topological matter field $\tilde{\bf G}$. These terms will depend on the topological field and the geometry of the space which will allow us to describe mixed states. First, we introduce in $\tilde{\bf G}$ a term proportional to |$\ket{\Phi}\bra{\Phi}$ |. Specifically we introduce a term $(m^{2}+\xi R)\ket{\Phi}\bra{\Phi}$ | where $R$ is the Ricci scalar and $\xi$ might include the case of conformal coupling $\xi=1/6$. This term can be also interpreted as a projector. Secondly, we introduce a term depending explicitly on the curvature of the manifold. Since $\tilde{\bf G}$ involves the metric for scalars, vectors and bivectors on equal footing, it is natural to consider a further term involving $\tilde{\bm{\mathcal{R}}}$ given by the direct sum of the Ricci scalar $R$, the Ricci tensor of elements $R_{\mu\nu}$ and the Riemann tensor of elements $R_{\mu\nu\rho\sigma}$, i.e.
$\tilde{\bm{\mathcal{R}}}$ $=$ $R\oplus\Big{(}R_{\mu\nu}dx^{\mu}\otimes dx^{\nu}\Big{)}$ (36)
@@
In the main body of this paper we will investigate only (bosonic) Dirac-Kähler matter fields. However gauge fields and fermionic Dirac fields can be included as well. For a discussion of the inclusion of Abelian gauge field see Appendix A.
Possibly this approach could be extended to include also higher-forms. However, for simplicity, we consider here only topological matter fields formed by the direct sum between a 0-form, a 1-form and a 2-form as this the minimal choice that will allow us to include in the action the Ricci scalar, the Ricci and the Riemann tensor explicitly.
@@ III.2 Entropic topological and geometrical action
We propose a statistical mechanics action formulated in terms of the quantum relative entropy between the metric $\tilde{g}$ and the metric $\tilde{\bf G}$ induced by the topological matter fields. In order to define our action we need to extend the notion of eigenvalues to metric tensors between bivectors. This will allow to define the entropy and the quantum relative entropy in this novel framework. For a detailed discussion of this mathematical background see Appendix B that constitutes also the reference for our notation. First of all we observe that the entropy associated to the metric $\tilde{g}$ remains zero. Indeed we define $\tilde{H}$ as
@@
$=$ $1\ln 1+\mbox{Tr}g\ln g^{-1}+\mbox{Tr}g_{(2)}\ln g_{(2)}^{-1}=0,$
where we have used Eq.(118) and the identity derived in the Appendix B (in Eq.(110)),
@@
Since the metric matrices $g_{(n)}$ have all their eigenvalues equal to one, using the notation developed in Appendix B we can express this Lagrangian as well as
@@
The resulting statistical mechanics action $\mathcal{S}$ associated to the Lagrangian $\mathcal{L}$ is strictly related to the Araki quantum relative entropy [37, 38, 17], and can be formally derived from a mathematical theory of quantum operators (see Appendix C). Specifically, the considered action is given by
@@
$\mathcal{L}$ $=$ $3{\beta}R-\alpha\bra{\Phi}{D}\tilde{g}^{-1}D\ket{\Phi}$ (46)
@@
$\bra{\Phi}{D}\tilde{g}^{-1}{D}\ket{\Phi}$ $=$ $|\nabla\phi|^{2}+|\nabla^{\mu}\omega_{\mu}|^{2}+\Big{|}\nabla^{\sigma}\bar{\omega}_{\sigma\mu}\Big{|}^{2}$ (47)